On a cloudy day, a raft of size 6x6 m floats on the surface of the pond that has a depth 0, 5 m and flat horizontal bottom. Determine the area of raft shadow on the bottom of the pond. Give the answer in [ m^2 ] and round it up to a whole. The refractive index of water is 4/3.
1
Expert's answer
2020-11-11T07:59:47-0500

"On a cloudy day" means that the light is diffused: the entire sky emits light; light rays have all possible directions. Thus, we have rays with maximum incidence angle of 90°. They will form the shadow.



From the right triangle we see that


"d=D-h\text{ tan}\alpha_2."

From optics, we know that


"\frac{\text{sin}\alpha_1}{\text{sin}\alpha_2}=n,\text{ since }\alpha_1=90°,\\\space\\
\text{sin}\alpha_2=\frac{1}{n},\\\space\\
\text{tan}\alpha_2=\frac{\text{sin}\alpha_2}{\sqrt{1-\text{sin}^2\alpha_2}}=\frac{1}{\sqrt{n^2-1}},\\\space\\
d=D-\frac{h}{\sqrt{n^2-1}}."

The area of the shadow is


"A=\frac{d^2}{2}=\frac{\bigg(D-\frac{0.5}{\sqrt{(n)^2-1}}\bigg)^2}{2}=\\\space\\
=\frac{\bigg(6\sqrt2-\frac{0.5}{\sqrt{(4/3)^2-1}}\bigg)^2}{2}=31\text{ m}^2."